We have productivity data from 2 mills

Data from Mill 1: 10,13,20,21,22,22,,25,28,29,32,34,38,40

Data from Mill 2: 10,10,12,15,18,18,18,19,19,21,22,22,23,23,24,27,28

  1. Calculate the mean and standard deviation for Mill 1
  2. Calculate the mean and standard deviation for Mill 2
  3. Assume the two are normal distributions, and the standard devisations are correct for the two populations the mills are drawn from. Finally assume the population mean for the two distributions are the same.
    1. Use the same variance rules to calculate the variance of the difference of the means.
    2. Calculate the Z score for the difference of the means.
    3. Determine the probability of seeing that Z score or one with a larger absolute value.

How do you find the answer to the question about determining the probability of seeing that z score or one with a larger absolute value? Please explain me the working or how you go about the answer.
Answer #1

- Answer: Given: Data from mill-.. 10,13,20,21,22,22,25,28,29) 36,38,40 Data from mill-2.3 10) 10)12, 15, 18,18,18,19,19,21,2

10 - 19.35 52) + 0-19:352)+ . +628-19.252 17-1 52 5.3143 10) p) Here we have to find v(&, -) v (q-2 ) = v(37) + v (83) 13 = 6

2p1272259) = 2 [1-P (222259)] > 2 [1-P6262:29] = 2 [1-20-9878] = 0.0244. cs Scanned with CamScanner